OpenStudy (anonymous):

Find the indicated trig value in the specified quadrant. Function: cot (x) = -3, quadrant IV, trig value: sin (x)

OpenStudy (jdoe0001):

what are the signs of the cosine and sine in the 4th Quadrant? are they both positive or negative or do they differ?

OpenStudy (anonymous):

Sine is negative, and cosine is positive.

OpenStudy (jdoe0001):

\(\bf cot(x) = -3\\ \quad \\ \quad \\ cot(x) = \cfrac{cos(x)}{sin(x)}\quad \textit{therefore}\quad cot(x) = -3\implies cot(x) = -\cfrac{3}{1}\\ \quad \\ \textit{since we're in the 4th Quadrant, sine is negative, thus}\\ \quad \\ cot(x) = -\cfrac{3}{1}\implies cot(x) = \cfrac{3}{-1}\implies \cfrac{a}{b}\\ \quad \\ \textit{to find "c" or the hypotenuse, we can use}\\ c^2 = a^2 + b^2\implies c = \sqrt{a^2+b^2}\\ \quad \\ \textit{keep in mind that }\qquad sin(x) = \cfrac{b}{c}\)